OpenStudy (anonymous):

The equation p=2.50v-500 gives the estimated profit "p" in dollars if v people visit the park In parts (a)-(b) find the answer by writing and solving an equation or inequality. a.For what number of visitors will the profit be about $2ooo b. 200 people visit. What is about the concession-stand profit for the day? c.For what number of visitors will the profit be at least $500 PLEASE HELP AND SHOW WORK :)

OpenStudy (agreene):

a. you need 2,000 profit so plug that in for p and solve for v: 2000=2.50v-500 b. it is saying v=200, so plug that in and evaluate for p. p=2.50(200)-500 c.since p must be at least 500 (but can be more) \[500\ge 2.50v-500\] solve for v.

OpenStudy (anonymous):

Thanks! :)